OpenStudy (anonymous):

Let f(x) = x^2 – 1 and g(x) = 1 – 2x. Find the indicated values. f(g(–1)) and g(f(–1))

OpenStudy (anonymous):

@misty1212 @Nnesha

OpenStudy (misty1212):

hey

OpenStudy (misty1212):

didn't we just do this?

OpenStudy (anonymous):

(different problem) same concept

OpenStudy (anonymous):

my aplogizes i copied the wrong one one sec

OpenStudy (misty1212):

no it isn't it is the same problem and the same concept

OpenStudy (anonymous):

f(g(–2)) and g(f(–2))

OpenStudy (anonymous):

(my aplogizes i copied the wrong one the above is the correct one)

OpenStudy (misty1212):

same thing for \[f(g(-2))\] find \(g(-2)\) first, then \(f\) of the result

OpenStudy (misty1212):

for \[g(f(-2))\]find \(f(-2)\) first then \(g\) of the result

OpenStudy (anonymous):

so g(-2)=1-2x f(-2)=x^2-1 ?

OpenStudy (anonymous):

so g=5 f=3

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

g(-2) =5 so you have to plug 5 into f(x)